Two books of a student are selected without random substitution where there are four examples of Statistics, three of Mathematics, and two of Finance. If X is the number of books of Statistics and Y is the number of volcanoes of selected Mathematics. construct a table that shows the values ​​of the joint probability distribution of X and Y.
Answer #1

Answer :

Number of books of Statistics = X

Number of volcanoes of selected Mathematics = Y

Total no.of books selected of student = 2

Examples of statistics = 4

Mathematics = 3

Finance = 3

therefore,

The table is as follows :

X
Y

0

1 2 3 Total
0 0.222 0.333 0.044 0

= 0.222 + 0.333 + 0.044 + 0

= 0.6

1

0.222

0.133 0 0

= 0.222 + 0.133

= 0.355

2 0.044 0 0 0 0.044
Total

= 0.222 + 0.222 + 0.044

= 0.489

= 0.333 + 0.133 + 0

= 0.467

0.044 0

= 0.488 + 0.466 + 0.044

= 1

Finally the no.of books can be selected in 2^1^0 ways